Write an equation of the line containing the given point and perpendicular to the given line.
(3,5);6x+y=2
The equation of the line is y= ____

Answer by nabla(475)

[1] Find the slope of the given line:
y=-6x+2
gives m=-6=slope
[2] A perpendicular line will have slope -1/m. In this case, -1/m=1/6
[3] Apply point slope form y-y1=m(x-x1) with point (3,5)=(x1,y1).
y-5=(1/6)(x-3)
y=(1/6)x-1/2+10/2
y=(1/6)x+9/2.
Finished